# QuaDreamer: Controllable Panoramic Video Generation for Quadruped Robots

### v1

@@ 2 Related Work
Diffusion Models for Conditional Generation. The diffusion model [40, 41, 42] was originally designed for image generation by learning the gradual denoising process from a Gaussian noise distribution to an image distribution. This paradigm achieves great success in high-quality video generation [43, 44, 43], such as SVD [44], AnimateDiff [45], and VideoCrafter [46]. Based on that, several works have introduced control signals and diverse control mechanisms to generate diverse controllable video, such as camera pose-driven video generation [47, 48, 49, 50], trajectory-based video generation [51, 52, 53, 54], and layout-to-image generation [55, 56, 14, 57]. Furthermore, works such as GeoDiffusion [58], TrackDiffusion [57], and MagicDrive [14] have shown that generated data can serve as datasets to boost downstream tasks. However, these works focus on global style and viewpoint transformations but fail to capture the subtle spatial jitters typical of quadruped robots. Therefore, this paper explores the potential of Image-to-Video (I2V) diffusion models in generating panoramic videos from the perspective of a quadruped robot, and, based on this, to advance the development of downstream perception models for quadruped robots.
@@ 5 Conclusion
We propose QuaDreamer, the first controllable panoramic data generation engine designed explicitly for quadruped robots. QuaDreamer is capable of producing highly controllable and realistic panoramic videos for perception tasks in complex environments. Limitations and future work. Due to the distortion and feature discontinuity in the boundary regions of panoramic images, the current model performs less effectively in boundary control compared to other areas. In the future, we plan to optimize the boundary regions by leveraging global information, thereby improving control performance. In addition, leveraging large language models to enable generation under coarsely annotated conditions also presents a valuable and underexplored avenue.

### v3

@@ 2 Related Work
Diffusion Models for Conditional Generation. The diffusion model [40, 41, 42] was originally designed for image generation by learning the gradual denoising process from a Gaussian noise distribution to an image distribution. This paradigm achieves great success in high-quality video generation [43, 44, 43], such as SVD [44], AnimateDiff [45], and VideoCrafter [46]. Based on that, several works have introduced control signals and diverse control mechanisms to generate diverse controllable videos, such as camera pose-driven video generation [47, 48, 49, 50], trajectory-based video generation [51, 52, 53, 54], and layout-to-image generation [55, 56, 14, 57]. Furthermore, works such as GeoDiffusion [58], TrackDiffusion [57], and MagicDrive [14] have shown that generated data can serve as datasets to boost downstream tasks. However, these works focus on global style and viewpoint transformations but fail to capture the subtle spatial jitters typical of quadruped robots. Therefore, this paper explores the potential of Image-to-Video (I2V) diffusion models in generating panoramic videos from the perspective of a quadruped robot, and, based on this, to advance the development of downstream perception models for quadruped robots.
@@ 5 Conclusion
We propose QuaDreamer, the first controllable panoramic data generation engine designed explicitly for quadruped robots. QuaDreamer explores a new data extraction method and is capable of producing highly controllable and realistic panoramic videos for perception tasks in complex environments.
6 Limitation
This paper focuses on the generation conditions under the small high-frequency jitter of quadruped robots, but lacks control generation for the movement of other degrees of freedom. In the future, we will explore other control methods for generation models. In addition, leveraging large language models to enable generation under coarsely annotated conditions also presents a valuable and underexplored avenue.